# Ramanujan’s function on small primes

### v1

@@ 1. Introduction
Lehmer asked whether Ramanujan’s tau function has any zeros, and he showed that, if it does, the smallest $n$ such that $\tau(n)=0$ is a prime number [Le47] . For this reason, as an example in a recent article [Bre25a] about matrix methods for the study of number-theoretic functions, we studied with computer calculations the function $n\mapsto\tau(p_{n})$ where $p_{1}=2$ and, more generally, $p_{n}$ denotes the $n^{th}$ prime. We noted that (within the range of our observations) the complex eigenvalues of certain matrices associated to this function exhibit rough regularities: the minimum modulus among the eigenvalues for the matrix associated to $\tau(p_{n})$ appears to oscillate in an approximately periodic fashion with $n$. For obvious reasons, we do not expect that these regularities are sustained without modification throughout the sequence $\{\tau(p_{n})\}$ or any of the other sequences we have examined.
2. Lemmas and definitions
To introduce the matrices that we will associate to the members of the sequence $\{\tau(p_{n})\}_{n\geq 1}$, we repeat some definitions and lemmas from our earlier article [Bre25a]. Although we supplied our own proofs of the lemmas (omitted here), they are well known in the theory of symmetric functions [M]. In the sequel, we denote the cardinality of a set $S$ by #S\#S and the determinant of a matrix $M$ by $|M|$.
@@ 2. Lemmas and definitions
With $f_{0}=1$, let $\overline{f}$ denote the sequence $f_{0},f_{1},f_{2},$.... Let $J_{n}(\overline{j})$ and $H_{n}(\overline{h})$ be the matrices
@@
We define deformations $J^{(c)}_{n}(\overline{j})$ of the $J_{n}(\overline{j})$ as the matrices
$\left(\begin{matrix}c&-1&0&\cdots&0\\ j_{1}&c&-2&\cdots&0\\ \vdots&\vdots&\vdots&\vdots&\vdots\\ j_{n-2}&j_{n-3}&\cdots&c&-n+1\\ j_{n-1}&j_{n-2}&j_{n-3}&\cdots&c\\ \end{matrix}\right).$
Remark 1.
At present we have no helpful insight into the theory of these deformed matrices except to say that (obviously) $J^{(c)}_{n}(\overline{j})=J_{n}\left(\overline{j}^{(c)}\right)$ where $\overline{j}^{(c)}=\{c,j_{1},j_{2},$...\}.
Now we state a corollary of the following theorem in Vein and Dale’s book  [Vein1999]. Let $A_{n}=$
@@
 [^1] Here we are merely codifying a method shown to us by J. López-Bonilla. (See Corollary 3.3 in our previous paper [Bre25a].)
@@
Let $C_{n}(x)=|(J_{n}(\overline{j})-xI_{n}|$, so that $C_{n}(x)=(-1)^{n}\chi({(J_{n}(\overline{j}))}(x)$. Then $C_{n}(x)=$
@@
3. Empirical observations about the sequence $\{\tau(p_{n})\}_{n\geq 1}$
3.1. Characteristic functions of the undeformed matrices representing the function $n\mapsto\tau(p_{n})$
Let $h_{\tau,p}(0)=1$ and $h_{\tau,p}(n)=\tau(p_{n})$ for positive integers $n$. Let $j_{\tau,p}(n)$ and $h_{\tau,p}(n)$ satisfy equation (D) together. Clearly, the values of $j_{\tau,p}(n)$ can be computed from equation (D1). Let $\Pi_{n}(x)=\chi(J_{n}(\overline{j_{\tau,p}}))=|xI-J_{n}(\overline{j_{\tau,p}})|$ be the characteristic polynomial of $J_{n}(\overline{j_{\tau,p}})$. It is clear that the following statements are equivalent: (a) $\tau(p_{n})=0$, (b) $|J_{n}(\overline{j_{\tau,p}})|=0$, and (c) the polynomial $\chi(J_{n}(\overline{j_{\tau,p}}))$ takes zero to zero.
In the range of our observations, the degree of $\Pi_{n}(x)$ is $n$. Writing $\Pi_{n}(x)=\sum_{k=0}^{n}a_{n,n-k}x^{n-k},$ [^2] See ‘undeformed tauprime2nov25no3.ipynb’ in the repository [Bre25].
(3) $a_{n,n-k}=(-1)^{k}k!\binom{n}{k}\tau(p_{k})$
(also in the range of our observations.) Equation (3) is immediate after Corollary 2.3.
@@ 3.1. Characteristic functions of the undeformed matrices representing the function $n\mapsto\tau(p_{n})$
But the number of such subsets $S$ is $\binom{n}{k}$, so the simplest way to reconcile Proposition 1 and equation (3) is to suppose that the structural symmetries of $J_{n}(\overline{j_{\tau,p}})$ somehow force $|J_{n}(\overline{j_{\tau,p}})[S,S]|=(-1)^{k}k!\tau(p_{k})$ just when #S=k\#S=k and $S\subseteq\{1,2,\ldots,n\}$. But this turns out to be too good to be true. [^3] See ‘too good.ipynb’in [Bre25]. Instead, there are (we hope) systematic cancellations coming from those symmetries. So far, we do not understand these cancellations.
3.2. Characteristic functions of the deformed matrices $J_{n}^{(c)}(\overline{j_{\tau,p}})$.
Let $\chi(J_{n}^{(1)}(\overline{j_{\tau,p}})=\Pi_{n}^{(1)}(x)$ (say.) As we said, in our earlier article [Bre25a] we described regularities in the zero sets of $\Pi_{n}^{(1)}(x)$. Specifically, let $V_{n}^{(1)}$ be the zeros of $\Pi_{n}^{(1)}(x)$ in $\mathbb{C}$ and let $\mu_{n}^{(1)}=\min_{v\in V_{n}^{(1)}}|v|$; then it appears that the graph of the pairs $(n,\mu_{n}^{(1)})$ form a nearly-periodic oscillating curve.
Figure 1. Minimum moduli of the roots of the $\chi_{\Pi^{(1)}_{n}}$
and their logarithms. (Deformed matrices with $c=1$.)
The deep spikes in the lower, logarithmic plot in Figure 1 suggest that the approach of the eigenvalues of $\Pi_{n}(x)$ (on which Lehmer’s question hang) to the origin may behave in a regular way, but the details are not yet understood by us because (among other reasons) our data set is very limited and because we do not know the details of the relationship of the eigenvalues of $\Pi_{n}(x)$ to those of the $\Pi_{n}^{(c)}(x)$. (The corresponding plots for $c=0$ exhibit similar behavior. [^4] See ‘deformed primetau 22oct25.ipynb’ in our repository [Bre25].)
The lower envelope of the lower plot of logarithms of minimum moduli for each $n$ in $\{2,3,...,400\}$ displayed in Figure 1 behaves regularly: the abscissas $n$ for a certain subset of the integers in $\{2,3,...,400\}$ of the points on the lower envelope ((corresponding to characteristic polynomial $\Pi_{n}(x)$) for the same subset), after the first three points are dropped, is
5 9 13 17 21 25 29 33 37
41 46 50 54 58 62 66 70 74
78 83 87 91 95 99 103 107 111
115 120 124 128 132 136 140 144 148
152 157 161 165 169 173 177 181 185
189 194 198 202 206 210 214 218 222
226 231 235 239 243 247 251 255 259
263 268 272 276 280 284 288 292 296
300 305 309 313 317 321 325 329 333
337 342 346 350 354 358 362 366 370
374 379 383 387 391 395.
The successive difference of these abscissas are
4 4 4 4 4 4 4 4 4
5 4 4 4 4 4 4 4 4
5 4 4 4 4 4 4 4 4
5 4 4 4 4 4 4 4 4
5 4 4 4 4 4 4 4 4
5 4 4 4 4 4 4 4 4
5 4 4 4 4 4 4 4 4
5 4 4 4 4 4 4 4 4
5 4 4 4 4 4 4 4 4
5 4 4 4 4 4 4 4 4
5 4 4 4.
where the value $5$ appears at the positions congruent to $1$ modulo $9$ (counting from $1$ but omitting $1$.)
Another description of this data is the following. When the abscissas $\{5,9,...,395\}$ are replaced by their residues modulo $4$, the result is
@@
2 3 3 3 3 3.
The size of the blocks of equal residues are $10$ for the first one and $9$ for each of the others. [^5] We are using data from the lower envelope of the logarithms of the minimum moduli. (The lower plot in Figure 1.) See “tauprime=h logarithmic envelopes prec100 n400 1dec25.ipynb” in [Bre25]. These observations suggest the residues of the sequence $\{\tau(p_{n})\}_{n=1,2,...}$ might be interesting to examine. (We have not done that yet.)
For contrast, we reproduce in Figure 2 another plot from our article [Bre25a]: the mimimum moduli for each $n$ for the undeformed characteristic polynomials $\Pi_{n}(x)$. We have not yet attempted the sort of analysis we just discussed for the $\Pi_{n}^{(1)}(x)$. The irregularities visible in Figure 2 do not seem promising to us in that regard. This is an obstacle to using these plots to look at Lehmer’s question, obviously, because the determinants of the matrices associated to the $\Pi_{n}^{(1)}(x)$ do not represent the map $n\mapsto\tau(p_{n})$ in a straightforward way. On the other hand, the issue may not be hopeless, in view of the way they were constructed from the values of that map.
Figure 2. Minimum moduli of the roots of undeformed $\chi_{\Pi_{n}}$
We studied the polynomials $\Pi_{n}^{(0)}(x))$. In the range n=1,2,,200n=1,2,...,200, they are monic of degree $n$, their $x^{n-1}$ coefficients are each zero, and they satisfy the following descriptions. [^6] See ‘c=0 deformed vs undeformed tauprime26oct25no3.ipynb’ in [Bre25]. Writing $\Pi_{n}^{(c)}(x)=\sum_{k=0}^{n}a_{c,n,k}x^{k}$, our observations satisfy the following equations:
(4) $a_{0,n,n-2}=-2^{3}*3*\binom{n}{2}$
(5) $a_{0,n,n-3}=2^{4}*3^{2}*\binom{n}{3}$
(6) $a_{0,n,n-4}=2^{2}*3^{3}*1061*\binom{n}{4}$
(7) $a_{0,n,n-5}=-2^{8}*3*16249*\binom{n}{5}$
(8) $a_{0,n,n-6}=2^{6}*3^{2}*5*13*32003*\binom{n}{6}$
(9) $a_{0,n,n-7}=-2^{6}*3^{4}*13*29*53*173*\binom{n}{7}$
(10) $a_{0,n,n-8}=-2^{4}*3^{3}*7*3181*2139307*\binom{n}{8}$
and [^7] See ’c=1&c=1vs0 deformed tauprime29oct25’ and ‘c=0,1 deformed tauprime2nov25.ipynb’ in [Bre25].
(11) $a_{1,n,n-2}=-23*\binom{n}{2}$
(12) $a_{1,n,n-3}=5*43*\binom{n}{3}$
(13) $a_{1,n,n-4}=7*16267*\binom{n}{4}$
(14) $a_{1,n,n-5}=-29*311*1447*\binom{n}{5}$
(15) $a_{1,n,n-6}=7*182111899*\binom{n}{6}$
(16) $a_{1,n,n-7}=-26573015431*\binom{n}{7}$
(17) $a_{1,n,n-8}=-5*53*17827*4318469*\binom{n}{8}$
In spite of appearances, the numerical coefficients of the binomial coefficients on the right side of the equations (11) - (17) are not always square-free. Apparently, these lists might be extended indefinitely.
Of course, the characteristic functions $\Pi_{n}^{(c)}(x)$ may also be written down in complete generality using Theorem 2.3.
The coefficients for $c=0$ and $c=1$ appear to satisfy another, related, description. [^8] See ‘c=0 deformed tauprime28oct25no3.ipynb’ in [Bre25]. For k=2,,200k=2,...,200, at least,
$a_{c,n,n-k}=a_{c,k-1,0}\cdot\binom{n}{k}=\left|J_{k-1}^{(c)}(\overline{j_{\tau,p}})\right|\cdot\binom{n}{k}$.

### v2

@@ 1. Introduction
Lehmer asked whether Ramanujan’s tau function has any zeros, and he showed that, if it does, the smallest $n$ such that $\tau(n)=0$ is a prime number [Le47] . For this reason, we made computer calculations about the function on positive integers
$\tau_{p}:n\mapsto\tau(p_{n})$ where $p_{n}$ denotes the $n^{th}$ prime. Using matrix machinery reproduced in Lemmas 2.1 and 2.2 below, within the range of our observations, the eigenvalues of the matrices $H_{n}$ and $J_{n}$ attached to this function by the two lemmas behave as follows: the minimum modulus among the eigenvalues for each of the two matrices associated to a particular $\tau_{p}(n)$ appear to oscillate with $n$. The connection to Lehmer’s question is that the associated matrix $H_{n}$ has a zero eigenvalue if and only the associated matrix $J_{n}$ does too, and, if they do, $\tau_{p}(n)=0$.
We show plots of the oscillations in Figure 1 and Figure 3 below. We also describe some pleasant patterns among the abscissas of the points on their lower envelopes that seem unlikely to persist unaltered: they seem to suggest that the effect of Ramanujan’s function on the prime sequence is to tune somehow the noise in it, and we do not know how that might work. We have looked at similar compositions of tau with other functions and, for them, this “tuning” is not so clear-cut.
2. Lemmas
In the context of the theory of symmetric functions, the lemmas in this section are well known. They appear for example, in MacDonald’s book [M]), but the proofs are not easy to locate. (In fact, we have not located them.) We have deposited home-made proofs of these statements in our depository [Bre25]111“Lemmas for ‘Ramanujan’s function on small primes’ ”. In the sequel, we denote the cardinality of a set $S$ by #S\#S and the determinant of a matrix $M$ by $|M|$.
@@ 2. Lemmas
With $f_{0}$ possibly defined as equal to one,, let $\overline{f}$ denote the sequence $\{f_{n}\}_{n\in\mathbb{Z}^{+}}$. Let $J_{n}(\overline{j})$ and $H_{n}(\overline{h})$ be the matrices
@@
Next we state a corollary of the following theorem in Vein and Dale’s book  [Vein1999].
Proposition 1.
(Vein and Dale, Theorem 4.23) Let $A_{n}=$
@@
 [^2] Here we are merely codifying an argument shown to us by J. López-Bonilla.
@@
Let $C_{n}(x)=|(J_{n}(\overline{j})-xI_{n}|$, so that $C_{n}(x)=(-1)^{n}\chi({(J_{n}(\overline{j}))}(x)$. By Proposition 1, $C_{n}(x)=$
@@
3. Empirical observations about the sequence $\{\tau_{p}(n)\}_{n\geq 1}$
3.1. Characteristic functions of the undeformed matrices of the form $J_{n}(\overline{j_{\tau_{p}}})$ representing the function $\tau_{p}$.
Let $h_{\tau_{p}}(0)=1$ and $h_{\tau_{p}}(n)=\tau_{p}(n)$ for positive integers $n$. Let $j_{\tau_{p}}(n)$ and $h_{\tau_{p}}(n)$ together satisfy equation (D). Given $h_{\tau_{p}}$, the elements of $\overline{j_{\tau_{p}}}=\{j_{\tau_{p}(n)}\}_{n=1,2,...}$ can be computed from equation (D1). By Lemma 2.2 (2), $\tau_{p}(n)=|J_{n}(\overline{j_{\tau_{p}}})|/n!$. Let $\Pi_{n}(x)=\chi(J_{n}(\overline{j_{\tau_{p}}}))(x)=|xI-J_{n}(\overline{j_{\tau_{p}}})|$ be the characteristic polynomial of $J_{n}(\overline{j_{\tau_{p}}})$. The following statements are clearly equivalent: (a) $\tau(p_{n})=0$,
(b) $|J_{n}(\overline{j_{\tau_{p}}})|=0$, and (c) $\Pi_{n}(0)=0$.
Writing [^3] See ‘undeformed tauprime2nov25no3.ipynb’ in the repository [Bre25]. $\Pi_{n}(x)=\sum_{k=0}^{n}a_{n,n-k}x^{n-k}$, Theorem 2.3 gives
(3) $a_{n,n-k}=(-1)^{k}k!\binom{n}{k}\tau_{p}(k).$
In (theoretical) practice, this result would allow us to bypass the construction of the matrices $J_{n}(\overline{j_{\tau_{p}}})$ in our computations, saving time and allowing us to extend our data sets. We have not found occasion to exploit this advantage; if a similar result applied to the $\chi(J_{n}^{(1)}(\overline{j_{\tau_{p}}}))(x)$, we would have used it; at present, this is not available to us.
@@ 3. Empirical observations about the sequence $\{\tau_{p}(n)\}_{n\geq 1}$
The number of terms in the sum is $\binom{n}{k}$, so the simplest way to reconcile Proposition 1 and equation (3) is to suppose that the structural symmetries of $J_{n}(\overline{j_{\tau_{p}}})$ somehow force $|J_{n}(\overline{j_{\tau_{p}}})[S,S]|=(-1)^{k}k!\tau_{p}(k)$ just when #S=k\#S=k and $S\subseteq\{1,2,\ldots,n\}$. But this turns out to be too good to be true. [^4] See ‘too good.ipynb’in [Bre25]. Instead, there must be systematic cancellations coming from those symmetries. So far, we have not understood these cancellations.
3.2. Characteristic functions of the deformed matrices $J_{n}^{(c)}(\overline{j_{\tau_{p}}})$.
[…]
Let $\chi(J_{n}^{(1)}(\overline{j_{\tau_{p}}}))=\Pi_{n}^{(1)}(x)$ (say.) (The corresponding objects for $c=0$ exhibit behaviors similar to those we go over here. [^5] See ‘deformed primetau 22oct25.ipynb’ in our repository [Bre25].) Let $V_{n}^{(1)}$ be the set of zeros of $\Pi_{n}^{(1)}(x)$ and let $\mu_{n}^{(1)}=\min_{v\in V_{n}^{(1)}}|v|$; then it appears that the graph of the pairs $(n,\mu_{n}^{(1)})$ lies on an oscillating curve.
Figure 1. Minimum moduli of the roots of the $\Pi^{(1)}_{n}(x)$
and their logarithms.
Figure 2. Minimum moduli of the roots of undeformed $\Pi_{n}(x)$.
The deep spikes in the lower, logarithmic plot in Figure 1 suggest that the approach of the roots of $\Pi_{n}^{(1)}(x)$ to the origin may behave in a predictable way, but the details are not clear to us because (among other reasons) our data is limited. Also, we do not know the details of the relationship of the roots of $\Pi_{n}(x)$ to those of the $\Pi_{n}^{(c)}(x)$, so the connection, if any, to Lehmer’s question is also unclear. Furthermore, we have no statement like equation (3) for the $\Pi_{n}^{(c)}(x)$, because we do not have an explicit description of the sequence $\{h_{n}\}$ represented in Lemma 2.2 by the matrices of which the $\Pi_{n}^{(c)}(x)$ are the characteristic polynomials. As we said above, such a description would permit us to sidestep the construction of the matrices we currently use to find these polynomials and the resulting time savings would in practice allow us to collect more data. More data would be welcome in view of the oscillations we are trying to check.
The lower envelope of the lower plot of logarithms of minimum moduli for each $n$ in $\{2,3,...,400\}$ displayed in Figure 1 appear to behave regularly: the abscissas of the points on the lower envelope form a set $\{5,9,...,395\}$, which, when they are replaced by their residues modulo $4$, make Table 1.
@@
2 3 3 3 3 3
Table 1. Residues modulo 4 of abscissas from the lower envelope of Figure 1.
The size of the blocks of equal residues are $10$ for the first one and $9$ for each of the others. [^6] We are using data from the lower envelope of the logarithms of the minimum moduli. (The lower plot in Figure 1.) See “tauprime=h logarithmic envelopes prec100 n400 1dec25.ipynb” in [Bre25].
For contrast, we reproduce in Figure 2 a plot of the mimimum moduli for each $n$ for the undeformed characteristic polynomials $\Pi_{n}(x)$. We have not yet attempted the sort of analysis we just discussed for the $\Pi_{n}^{(1)}(x)$. The irregularities visible in Figure 2 do not seem promising to us in that regard. This is an obstacle to using these plots to look at Lehmer’s question because the determinants of the matrices associated to the $\Pi_{n}^{(1)}(x)$ do not represent the function $\tau_{p}$ in a straightforward way.
3.3. Characteristic functions of deformed matrices $H_{n}^{(1)}(\overline{h_{\tau_{p}}})$.
 [^7] See our notebook “tauprime=j H matrix4dec25no2.ipynb” in [Bre25].
Given the sequence $\{\tau_{p}(n)\}_{n=1,2,...}=$ (say) $\{j_{\tau_{p}}(n)\}_{n=1,2,...}=\overline{j_{\tau_{p}}}$, we
construct a sequence $\overline{h_{\tau_{p}}}=\{h_{\tau_{p}}(n)\}_{n=1,2,...}$ by solving the convolution (D1). By Lemma 2.2, $\tau_{p}(n)=(-1)^{n+1}|H_{n}(\overline{h})|$. Then we construct matrices
[…]
Figure 3. Minimum moduli and their logarithms for zeros of the $\Psi^{(1)}_{n}(x)$.
The table of modulo $4$ residues for the abscissas of the lower envelopes of the plots in Figure 3 is not as regular as the corresponding table for Figure 1; the lengths of the strings of constant residue vary:
0 0 2 2 3 3 3 3 3
0 0 0 0 0 1 1 1 1
1 1 2 2 2 2 2 3 3
3 3 3 3 0 0 0 0 0
1 1 1 1 1 1 2 2 2
2 2 2 2 3 3 3
Table 2. Residues modulo 4 of abscissas from thelower envelope of Figure 3.